(Solve only question 1) A $6M investment is considered by an electric bike manufacturing company to add a new production line for its new product, electric skateboards. The company has commissioned an exploratory study of where to place the new production line and which type of equipment to use. There are three types of machines to choose from for the company to install on the new assembly line. The machines have zero salvage value at the end of 10-year planning horizon. The company must select at least two alternatives from (i) Loan, (ii) Common Stocks, (iii) Preferred Stocks, and (iv) Retained Earnings to obtain the required amount of capital for the investment. Each of these capital sources could provide $3M to support the project. The company is anticipating rapid product penetration and aggressive growth after addition of the new production line.

Options 1 Capital Sources Description 1 Loan Interest Rate = 8 %, Compounded Semi Annually Payback Method: Plan 3

2 Preferred Stock Dividend = $7, Price = $100, Brokerage Fee = $1

3 Common Stock Dividend = $5, Price = $100, Growth Rate = 4%

4 Retained Earning

Question 1) Only chose Loan & Common stock from the above option and form the capital for the investment,& calculate WACC.

Question 2) Calculate MARR
Answer #1

Cost of Loan = Kd = Effective Interest Rate = 8% compunded semi annually = 8.16% (100+4%+4%-100=8.16)

Cost of Equity = Ke = Dividend/Price + Growth (Gordon's Model)

= 5/100% + 4% = 9%

Calculation of WACC :

Source of fund Amount Cost Weight Weighted Cost
Loan 3M 8.16% 50% 4.08%
Common Stock 3M 9% 50% 4.50%
Total 6M 8.58%

Thus, WACC = 8.58% p.a.